OpenStudy (anonymous):

What point in the feasible region maximizes the objective function. I just need help graphing.

OpenStudy (anonymous):



hero (hero):

Have you tried graphing it using desmos.com?

OpenStudy (anonymous):

I tried using the desmos calc and I don't think I put it in right. I put in all the info and the feasible region has only 3 points and it doesn't look correct.

hero (hero):

You do realize that (0,0) is a point as well, right?

OpenStudy (anonymous):

see, idk how you do that. I got (-3,0) (1.5,1.5) (3,0) and then you said (0,0) which didn't look like it was a point on mine

hero (hero):

For this case, I assume this is what you have: https://www.desmos.com/calculator/sb8qy6dhpz

OpenStudy (anonymous):

Do you just copy and paste the website? So I can show you what I have

hero (hero):

The inequality y <= 1/3(x) + 3 doesn't even factor into it because it only intersects with a point on the edge of the region. So you truly do only have three points.

hero (hero):

Take a look at that last link I posted

OpenStudy (anonymous):

I did and mine looks nothing like that haha

hero (hero):

Hmmm

hero (hero):

Well, that's the correct region, and it only has three points like you were saying.

OpenStudy (anonymous):

how can I show you what I have

hero (hero):

You would have to sign in with google to be able to generate links

OpenStudy (anonymous):

https://www.desmos.com/calculator/6184og09zb

hero (hero):

Oh the numbers are different from what I thought. Hang on

OpenStudy (anonymous):

So, the points you gave me earlier in one of the links are not correct?

hero (hero):

No, forget those other links. I'll post the the right graph this time. If you want this to work for you, you have to learn how to plot line segments

hero (hero):

There we go

OpenStudy (anonymous):

alright, although I have no idea how you plotted the line segments.

hero (hero):

Basically what you see is four sets of things plotted. First the equations: x = 0 y = 0 y = (1/3)x + 1 3 = y + x Then I plot the intersecting points of those Then I plot the line segments. Study the last 4 segments. They correspond with the graph. The stuff between the brackets determine the length of each segment along a given line with respect to the x or y axis.

hero (hero):

Study the brackets

OpenStudy (anonymous):

That will take a while for me to figure out. I get the point of it and how they relate but if I were to try and do it myself I would have some trouble. the points are: (0,0) (0,1) (3,0) (1.5,1.5) right?

hero (hero):

Yes, those are the points of interest.

OpenStudy (anonymous):

Ill keep it open though so I can go over it more tomorrow

hero (hero):

If you signed into google you can save the graph on your desmos profile, then access it later.

OpenStudy (anonymous):

alright, If I tell you the objective function and what I got for the answer will you check it?

hero (hero):

Okay, yeah real quick.

OpenStudy (anonymous):

I saved it. and k, hold on let me finish. I am almost done

OpenStudy (anonymous):

C=5x-4y

OpenStudy (anonymous):

(0,0): 5(0)-4(0)=0 (0,1): 5(0)-4(1)=-4 (3,0): 5(3)-4(0)=15 (1.5,1.5): 5(1.5)-4(1.5)=1.5 So the solution is (3,0) 15

hero (hero):

Correct

OpenStudy (anonymous):

alright, thank you for helping.